# JHU Applied Physics Laboratory, 11100 Johns Hopkins Rd, Laurel, MD 20723, USA

### v1

@@ 6 Discussion
When giant planets are included, it is possible to extend the probed sample to equilibrium temperatures of 2500 K. This is done by Fu et al. (2017) and Edwards et al. (2023), who find an overall increasing feature size with equilibrium temperature. They also see a local maximum in feature size at approximately (see Fig. 12), aligned with the featureless spectrum of HD 86226 c. This discrepancy might be caused by the differences between giant planet and sub-Neptune atmospheres. While giant planets in general have approximately solar-metallicity atmospheres, recent JWST observations suggest that many sub-Neptunes have metal-rich atmospheres (e.g., GJ 1214 b, GJ 9827 d, TOI-270 d, and GJ 3090 b Kempton et al., 2023; Roy et al., 2023; Benneke et al., 2024; Ahrer et al., 2025). High-metallicity atmospheres have a smaller scale height, which decreases the spectral feature amplitude. In addition, their reservoir of potentially cloud-forming metals is larger. This may allow for an increased formation of clouds on sub-Neptunes, further muting their spectral features. For giant planets, Gao et al. (2020) explain the local feature size maximum with a sinking of silicate clouds to higher atmospheric pressures as the equilibrium temperature of a planet decreases. In sub-Neptunes, different atmospheric metallicities and temperature structures may prevent the clouds from sinking.
@@ 6.1 The hazy scenario
Since CO and CO2 are possible haze precursors, our haze models based on CH4, HCN, and C2H2 serve as a calculation with the lower limit of the haze production rate. In this lower limit, the haze produced is not sufficient to mute the atmospheric features of HD 86226 c. If aerosols shape the muted spectrum of HD 86226 c, these are either hazes primarily formed through the photodissociation of CO and CO2, or cloud species.
@@ 6.2 The cloudy scenario
Many cloud species may contribute to opaque high-temperature atmospheres, as shown in Fig. 8. We particularly focus on MnS, MgSiO3, and Fe clouds, as the corresponding condensation curves indicate that these species could condense at the atmospheric pressures that are probed by our observations, . The presence of other silicate species like Mg2SiO4 and SiO2 (see e.g., Grant et al., 2023; Dyrek et al., 2024; Inglis et al., 2024) is also plausible for this planet, although these would condense at higher atmospheric pressures than MgSiO3 and are likely to be less abundant (Lodders & Fegley Jr, 2006; Visscher et al., 2010). Most cloud species typically predicted on smaller Neptune-sized planets, such as Na2S, KCl, and ZnS, evaporate at the temperatures of HD 86226 c.
@@
Another point to consider is that not all of these particles may be available for cloud droplet formation. Some may participate in other chemical processes or be inhibited by rapid vertical mixing. Vertical mixing can also replenish certain species, thereby sustaining the gas-phase reservoir available for condensation. In this case, the necessary species may be supplied from deeper atmospheric layers. In contrast, larger droplets may also sink to deeper layers as a consequence of gravitational settling. Our cloud prescription applies a simplifying assumption that clouds are homogenously distributed above altitudes where the condensation curves cross the pressure-temperature profile. In the presence of large-scale zonal flows like the ones present in hot Jupiters (see e.g., Fortney et al., 2021), this description may be inaccurate. These flows may mix condensates below the pressure level predicted by this crossing. Furthermore, they would also preferably transport a specific size range of condensate droplets. Larger droplets would sink to lower atmospheric layers, while smaller droplets could be lofted. Depending on the position and extent of such a flow in the planet’s atmosphere, a vertically homogenous cloud droplet size distribution does not reproduce this effect.
@@ 6.3 The metal-rich scenario
Recent JWST observations of the sub-Neptunes suggest that metal-rich atmospheres are common on sub-Neptunes (e.g., GJ 1214 b, GJ 9827 d, TOI-270 d, and GJ 3090 b Kempton et al., 2023; Roy et al., 2023; Benneke et al., 2024; Ahrer et al., 2025). The featureless spectrum of HD 86226 c can also be explained by a high mean molecular weight atmosphere alone, without the need for aerosols.
@@
A high mean molecular weight atmosphere of HD 86226 c in combination with the presence of the outer giant companion HD 86226 b (Teske et al., 2020) would imply interesting scenarios for its formation history. The high mean molecular weight atmosphere could originate from the accretion of water vapor from an enriched inner disc (Bitsch et al., 2021), which might have formed from inward drifting and evaporating water pebbles (Schneider & Bitsch, 2021). However, the outer growing giant beyond the water ice line could open a gap in the disc to block the inward-flowing pebbles. This would reduce the vapor enrichment in the inner regions of the disc and thus reduce the mean molecular weight of a planet forming in it. However, a high molecular weight atmosphere of HD 86226 c would indicate two possible scenarios: The giant planet could have formed interior to the water ice line of the system, where the water ice pebbles have already evaporated. Alternatively, HD 86226 c could have formed exterior to the ice line and migrated inwards, before the outer giant planet could block the inward flow of pebbles (e.g., Bitsch et al., 2021). This system is the second system after HAT-P-11, where a formation analysis like this is possible (see Chatziastros et al., 2024). Knowing the composition of the inner planets can therefore help us to learn about the different formation scenarios of outer giant planets, a puzzle also present in the Solar System (e.g., Kruijer et al., 2017).
@@ 6.4 Possible interior structures of HD 86226 c
The combination of mass, radius, and atmospheric metallicity alone presents a degenerate problem in constraining the size and the composition of the deep core (Kramm et al., 2011; Otegi et al., 2020). Therefore, the core composition could range from pure rock to a mixture of rock and iron with dissolved volatiles (Vazan et al., 2022; Luo et al., 2024). Determining the atmospheric abundances of H2O, CO2, and CH4 with future JWST transmission spectra may help constrain the composition of the deep interior, particularly the water-to-hydrogen ratio (Yang & Hu, 2024). However, further modeling is required to confirm whether this is also applicable to hot (1000 K) sub-Neptunes like HD 86226 c.
@@ 6.5 How to solve the cloud-metallicity degeneracy
This degeneracy could be broken by observing HD 86226 c with the NIRSpec/G395H instrument of JWST, as CO2 has a prominent feature above 3 m that is highly sensitive to the atmospheric metallicity. Figure 13 displays the predicted spectra for different possible atmospheric edge-cases on HD 86226 c, as derived from the HST data: high-metallicity () atmospheres with a grey cloud-deck below bar or above bar, and a low-metallicity () atmosphere with a cloud-deck above bar. For each of these scenarios, we used pRT to generate spectra from all retrieval posteriors (see Sect. 5.1) that match the respective description. For these spectra, we included the opacities of H2O, CO, and CO2, as these are the dominant species predicted for planets like HD 86226 c by Moses et al. (2013). We used the median of these spectra to obtain a model spectrum for each scenario, respectively. For each of these three model spectra, we used pandexo (Batalha et al., 2017) to simulate NIRSpec/G395H data of three transit observations. On the simulated data, we additionally added a random noise.
@@
To probe the clouds directly, JWST/MIRI observations may enable us to discern better which cloud species are muting the spectra of HD 86226 c. If silicate clouds are present in HD 86226 c, it could be possible to observe the silicate features around (see Fig. 9 and Marley & Robinson, 2015). However, directly detecting silicate features in the possibly metal-enriched atmosphere of HD 86226 c may require much observing time, as previously simulated by Piette et al. (2023). At shorter wavelengths, Rayleigh scattering on small cloud particles might produce a spectral slope. High sensitivity observations at visual wavelengths could therefore provide a further probe of the cloud coverage on HD 86226 c. Finally, another possibility for detecting and characterizing clouds on the planet is to measure the occultation depth, or phase curve, of the planet (see e.g., Kempton et al., 2023; Coulombe et al., 2025). In the presence of silicate clouds, the planet is expected to have a high albedo, reflecting a large fraction of the incoming radiation (e.g., Garcia Munoz & Isaak, 2015).
@@ 7 Summary
For the HST/WFC3 data reduction, we used the open-source PACMAN and Eureka! packages. In contrast to typical WFC3 spatial scanning observations, the high scan rate for this bright star resulted in shifts in the position of the spectral trace on the detector. We observed shifts of up to two pixels, resulting in a flux measurement that was correlated with the position of the trace. To account for this effect, we implemented a function in PACMAN that decorrelates flux and trace position for both forward and reverse scan directions. For the data reduction with Eureka!, we modified the pipeline to allow for the simultaneous fitting of multiple HST transits. We find that PACMAN and Eureka! result in qualitatively similar spectra, but note that Eureka! does not correct the trace position instabilities, resulting in a higher level of correlated noise in the light curves.
The optical photometric monitoring reveals HD 86226 as a quiet host star with a standard deviation in nightly variability of in the V band. Our nine transit observations do not show evidence of stellar flares or spot-crossings of the planet.
The transmission spectrum of HD 86266 c is featureless and consistent within with a constant transit depth of ppm. This corresponds to an amplitude of just scale heights for an H/He-dominated atmosphere. Based on an atmospheric retrieval analysis with petitRADTRANS, we rule out a cloud-free, solar metallicity atmosphere with a confidence of . However, the existing data cannot unambiguously distinguish whether aerosols or a high-metallicity atmosphere are responsible for the muted spectral features. In the absence of clouds, the observed spectrum could be produced by a metal-enriched atmosphere with ( confidence lower limit). In addition to the atmospheric retrieval analysis, we also compare the spectrum to predictions from forward models of clouds and hazes. In contrast to most sub-Neptunes, we find that the high temperature of HD 86226 c prohibits the formation of organic hazes with CH4, HCN, and C2H2 as haze precursors. On the other hand, it allows for the formation of refractory cloud species, such as MnS, MgSiO3, and Fe clouds, similar to those expected for hot Jupiters. All of these cloud species could plausibly match the observations.
The featureless spectrum of the hot and small sub-Neptune HD 86226 c is surprising in the context of previously observed trends for sub-Neptune atmospheres, which found that hotter planets should have larger amplitude spectral features (e.g., Brande et al., 2024). Follow-up observations with JWST may clarify the physical processes and atmospheric chemistry that shape the observed spectrum. Determining this planet’s exact atmospheric composition will show if it hosts a cloud species atypical for sub-Neptunes or if it was formed with a metal-enriched atmosphere.
@@ Appendix A PACMAN light curve fits
For the spectroscopic bins, the increased RMS makes the deviation from a linear flux baseline undetectable, which is why the RMS behaves like white noise in most wavelength bins. Minor deviations from this are seen in the bin at , as there is an artifact on the WFC3 detector. Furthermore, bins at the long-wavelength edge show slight signs of red noise. Since this contribution is minor, we did not account for it further.

### v2

@@ 6 Discussion
When giant planets are included, it is possible to extend the probed sample to equilibrium temperatures of 2500 K. This is done by Fu et al. (2017) and Edwards et al. (2023), who find an overall increasing feature size with equilibrium temperature. They also see a local maximum in feature size at approximately (see Fig. 12), aligned with the featureless spectrum of HD 86226 c. This discrepancy might be caused by the differences between giant planet and sub-Neptune atmospheres. While giant planets in general have approximately solar-metallicity atmospheres, recent JWST observations suggest that many sub-Neptunes have metal-rich atmospheres (e.g., GJ 1214 b, GJ 9827 d, TOI-270 d, and GJ 3090 b Kempton et al., 2023; Roy et al., 2023; Piaulet-Ghorayeb et al., 2024; Benneke et al., 2024; Ahrer et al., 2025). High-metallicity atmospheres have a smaller scale height, which decreases the spectral feature amplitude. In addition, their reservoir of potentially cloud-forming metals is larger. This may allow for an increased formation of clouds on sub-Neptunes, further muting their spectral features. For giant planets, Gao et al. (2020) explain the local feature size maximum with a sinking of silicate clouds to higher atmospheric pressures as the equilibrium temperature of a planet decreases. In sub-Neptunes, different atmospheric metallicities and temperature structures may prevent the clouds from sinking.
@@ 6.1 The hazy scenario
Since CO and CO2 are possible haze precursors, our haze models based on CH4, HCN, and C2H2 serve as a calculation with the lower limit of the haze production rate. In this lower limit, the haze produced is not sufficient to mute the atmospheric features of HD 86226 c. If aerosols shape the muted spectrum of HD 86226 c, these are either hazes primarily formed through the photodissociation of CO and CO2 or cloud species.
@@ 6.2 The cloudy scenario
Many cloud species may contribute to opaque high-temperature atmospheres, as shown in Fig. 8. We particularly focus on MnS, MgSiO3, and Fe clouds, as the corresponding condensation curves indicate that these species could condense at the atmospheric pressures that are probed by our observations, . The presence of other silicate species such as Mg2SiO4 and SiO2 (see e.g., Grant et al., 2023; Dyrek et al., 2024; Inglis et al., 2024) is also plausible for this planet, although these would condense at higher atmospheric pressures than MgSiO3 and are likely to be less abundant (Lodders & Fegley Jr, 2006; Visscher et al., 2010). Most cloud species typically predicted on smaller Neptune-sized planets, such as Na2S, KCl, and ZnS, evaporate at the temperatures of HD 86226 c.
@@
Another point to consider is that not all of these particles may be available for cloud droplet formation. Some may participate in other chemical processes or be inhibited by rapid vertical mixing. Vertical mixing can also replenish certain species, thereby sustaining the gas-phase reservoir available for condensation. In this case, the necessary species may be supplied from deeper atmospheric layers. In contrast, larger droplets may also sink to deeper layers as a consequence of gravitational settling. Our cloud prescription applies a simplifying assumption that clouds are homogenously distributed above altitudes where the condensation curves cross the pressure-temperature profile. In the presence of large-scale zonal flows such as the ones present in hot Jupiters (see e.g., Fortney et al., 2021), this description may be inaccurate. These flows may mix condensates below the pressure level predicted by this crossing. Furthermore, they would also preferably transport a specific size range of condensate droplets. Larger droplets would sink to lower atmospheric layers, while smaller droplets could be lofted. Depending on the position and extent of such a flow in the planet’s atmosphere, a vertically homogenous cloud droplet size distribution does not reproduce this effect.
@@ 6.3 The metal-rich scenario
Recent JWST observations of the sub-Neptunes suggest that metal-rich atmospheres are common on sub-Neptunes (e.g., GJ 1214 b, GJ 9827 d, TOI-270 d, and GJ 3090 b Kempton et al., 2023; Roy et al., 2023; Piaulet-Ghorayeb et al., 2024; Benneke et al., 2024; Ahrer et al., 2025). The featureless spectrum of HD 86226 c can also be explained by a high mean molecular weight atmosphere alone, without the need for aerosols.
@@
A high mean molecular weight atmosphere of HD 86226 c in combination with the presence of the outer giant companion HD 86226 b (Teske et al., 2020) would imply interesting scenarios for its formation history. The high mean molecular weight atmosphere could originate from the accretion of water vapor from an enriched inner disk (Bitsch et al., 2021), which might have formed from inward drifting and evaporating water pebbles (Schneider & Bitsch, 2021). However, the outer growing giant beyond the water ice line could open a gap in the disk to block the inward-flowing pebbles. This would reduce the vapor enrichment in the inner regions of the disk and thus reduce the mean molecular weight of a planet forming in it. However, a high molecular weight atmosphere of HD 86226 c would indicate two possible scenarios: The giant planet could have formed interior to the water ice line of the system, where the water ice pebbles have already evaporated. Alternatively, HD 86226 c could have formed exterior to the ice line and migrated inward, before the outer giant planet could block the inward flow of pebbles (e.g., Bitsch et al., 2021). This system is the second system after HAT-P-11, where such a formation analysis is possible (see Chatziastros et al., 2024). Knowing the composition of the inner planets can therefore help us learn about the different formation scenarios of outer giant planets, a puzzle also present in the Solar System (e.g., Kruijer et al., 2017).
@@ 6.4 Possible interior structures of HD 86226 c
The combination of mass, radius, and atmospheric metallicity alone presents a degenerate problem in constraining the size and the composition of the deep core (Kramm et al., 2011; Otegi et al., 2020). Therefore, the core composition could range from pure rock to a mixture of rock and iron with dissolved volatiles (Vazan et al., 2022; Luo et al., 2024). Determining the atmospheric abundances of H2O, CO2, and CH4 with future JWST transmission spectra may help constrain the composition of the deep interior, particularly the water-to-hydrogen ratio (Yang & Hu, 2024). However, further modeling is required to confirm whether this is also applicable to hot (1000 K) sub-Neptunes similar to HD 86226 c.
@@ 6.5 How to solve the cloud-metallicity degeneracy
This degeneracy could be broken by observing HD 86226 c with the NIRSpec/G395H instrument of JWST, as CO2 has a prominent feature above 3 m that is highly sensitive to the atmospheric metallicity. Figure 13 displays the predicted spectra for different possible atmospheric edge-cases on HD 86226 c, as derived from the HST data: high-metallicity () atmospheres with a gray cloud-deck below bar or above bar, and a low-metallicity () atmosphere with a cloud-deck above bar. For each of these scenarios, we used pRT to generate spectra from all retrieval posteriors (see Sect. 5.1) that match the respective description. For these spectra, we included the opacities of H2O, CO, and CO2, as these are the dominant species predicted for planets similar to HD 86226 c by Moses et al. (2013). We used the median of these spectra to obtain a model spectrum for each scenario, respectively. For each of these three model spectra, we used pandexo (Batalha et al., 2017) to simulate NIRSpec/G395H data of three transit observations. On the simulated data, we additionally added a random noise.
@@
Observations with JWST/MIRI may enable a better distinction of the cloud species that mute the spectra of HD 86226 c, by probing the clouds directly. If silicate clouds are present in HD 86226 c, it may be possible to observe the silicate features around (see Fig. 9 and Marley & Robinson, 2015). However, directly detecting silicate features in the possibly metal-enriched atmosphere of HD 86226 c may require much observing time, as previously simulated by Piette et al. (2023). At shorter wavelengths, Rayleigh scattering on small cloud particles might produce a spectral slope. High sensitivity observations at visual wavelengths could therefore provide a further probe of the cloud coverage on HD 86226 c. Finally, another possibility for detecting and characterizing clouds on the planet is to measure the occultation depth, or phase curve, of the planet (see e.g., Kempton et al., 2023; Coulombe et al., 2025). In the presence of silicate clouds, the planet is expected to have a high albedo, reflecting a large fraction of the incoming radiation (e.g., Garcia Munoz & Isaak, 2015).
@@ 7 Summary
For the HST/WFC3 data reduction, we used the open-source PACMAN and Eureka! packages. In contrast to typical WFC3 spatial scanning observations, the high scan rate for this bright star resulted in shifts in the position of the spectral trace on the detector. We observed shifts of up to two pixels, resulting in a flux measurement that was correlated with the position of the trace. To account for this effect, we implemented a function in PACMAN that decorrelates flux and trace position for both forward and reverse scan directions. For the data reduction with Eureka!, we modified the pipeline to allow for the simultaneous fitting of multiple HST transits. We find that PACMAN and Eureka! result in qualitatively similar spectra, but we note that Eureka! does not correct the trace position instabilities, resulting in a higher level of correlated noise in the light curves.
The optical photometric monitoring revealed HD 86226 as a quiet host star with a standard deviation in nightly variability of in the V band. Our nine transit observations do not show evidence of stellar flares or spot crossings of the planet.
The transmission spectrum of HD 86266 c is featureless and consistent within with a constant transit depth of ppm. This corresponds to an amplitude of just scale heights for an H/He-dominated atmosphere. Based on an atmospheric retrieval analysis with petitRADTRANS, we ruled out a cloud-free solar metallicity atmosphere with a confidence of . However, the existing data cannot unambiguously distinguish whether aerosols or a high-metallicity atmosphere are responsible for the muted spectral features. In the absence of clouds, the observed spectrum could be produced by a metal-enriched atmosphere with ( confidence lower limit). In addition to the atmospheric retrieval analysis, we also compared the spectrum to predictions from forward models of clouds and hazes. In contrast to most sub-Neptunes, we find that the high temperature of HD 86226 c prohibits the formation of organic hazes with CH4, HCN, and C2H2 as haze precursors. On the other hand, it allows for the formation of refractory cloud species such as MnS, MgSiO3, and Fe clouds, which are similar to those expected for hot Jupiters. All of these cloud species could plausibly match the observations.
The featureless spectrum of the hot and small sub-Neptune HD 86226 c is surprising in the context of previously observed trends for sub-Neptune atmospheres, namely that hotter planets should have larger amplitude spectral features (e.g., Brande et al., 2024). Follow-up observations with JWST may clarify the physical processes and atmospheric chemistry that shape the observed spectrum. Determining this planet’s exact atmospheric composition will show if it hosts a cloud species atypical for sub-Neptunes or if it was formed with a metal-enriched atmosphere.
@@ Acknowledgements.
- Blecic et al. (2025) Blecic, J., Baker, A., Dobbs-Dixon, I., et al. 2025, manuscript submitted to The Astrophysical Journale
@@
- Piaulet-Ghorayeb et al. (2024) Piaulet-Ghorayeb, C., Benneke, B., Radica, M., et al. 2024, ApJ, 974, L10
@@ Appendix A PACMAN light curve fits
For the spectroscopic bins, the increased RMS makes the deviation from a linear flux baseline undetectable, which is why the RMS behaves similar to white noise in most wavelength bins. Minor deviations from this are seen in the bin at , as there is an artifact on the WFC3 detector. Furthermore, bins at the long-wavelength edge show slight signs of red noise. Since this contribution is minor, we did not account for it further.